Evaluate the integration by substitution ∫▒sin2θ/√(〖cos〗^2+16) dθ


1
Expert's answer
2021-04-29T17:21:57-0400

We have to find value of the integral


I = "\int \dfrac{sin2 \theta}{\sqrt{cos^2{\theta}+16}}d\theta"


Using the substitution,


"sin2\theta = 2cos\theta sin\theta"

= "\int- \dfrac{-1}{\sqrt{cos^2{\theta}+16}}2sin\theta cos\theta d\theta"


Substitute "u = cos^2{\theta}+16"

"\dfrac{du}{d\theta} = -2cos\theta sin\theta"

Hence,

"d\theta = -\dfrac{du}{2cos\theta sin\theta}"

Hence integral will become


"-\int \dfrac{du}{\sqrt{u}}"

"= -2\sqrt{u}"

Using substitution "u = cos^2{\theta}+16"


We get , "I = -2\sqrt{cos^2{\theta}+16} + C"